When 0.0167 g of a compound with a heat of combustion of -41.6 kJ/mole is burned
in a calorimeter, 28.15 J of heat is released. What is the molar mass of the compound?
1
Expert's answer
2021-10-15T02:58:47-0400

∆H"^{\theta}{_c}=-41.6kJmol^{-1}"

"=-41600Jmol^{-1}"

Number of moles that releases 28.15J of heat

"=\frac{28.15}{41600} moles"


Therefore "\frac{28.15}{41600} moles" is equivalent to


0.0167g


Molar mass "=0.0167÷(\frac{28.15}{41600})"


"=0.0167×\frac{41600}{28.15}"


= 24.6792g